OpenStudy (anonymous):

please help. 8 questions, 10 minutes and will fan any help and medal for best response. im running out of time please help, please help.

OpenStudy (anonymous):

Which word completely describes the polygon? A.equilateral B.equiangular C.regular

OpenStudy (texaschic101):

equiangular because all the angles are the same

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

hold on and ill put another question up.

OpenStudy (anonymous):

The length of a rectangle is 15 feet. The width is 8 feet. What is the perimeter of the rectangle? A.46 feet B.23 feet C.38 feet D.120 feet

OpenStudy (texaschic101):

P = 2L + 2W P = 2(15) + 2(8) P = 30 + 16 P = 46

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

What is the area of the trapezoid? A.1344 square inches B.168 square inches C.132 square inches D.96 square inches

OpenStudy (texaschic101):

A = (a + b) / 2 * h A = (14 + 8)/2 * 12 A = 22/2 * 12 A = 11 * 12 A = 132 sq inches

OpenStudy (anonymous):

thx. i really get confused on these type of things. thanks so much for ur help. i have 3 more questions if u could help.

OpenStudy (texaschic101):

I can try...they can be confusing. You just have to know the formulas

OpenStudy (anonymous):

A rectangle has an area of 432 square centimeters and a width of 18 centimeters. What is the length of the rectangle? A.18 centimeters B.36 centimeters C.24 centimeters D.48 centimeters

OpenStudy (texaschic101):

A = L * W 432 = 18L -- divide by 18 432/18 = L 24 = L

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the area of the rectangle? A.35 square feet B.17.5 square feet C.78 square feet D.76 square feet

OpenStudy (texaschic101):

A = L * W A = 8 * 9.5 A = 76

OpenStudy (anonymous):

Use graph paper and a straightedge to draw the figure. The set of points (–1, 1), (5, 1), (–3, –2), and (3, –2) identifies the vertices of a quadrilateral. What is the most specific description of the figure drawn? A.parallelogram B.rectangle C.trapezoid D.square

OpenStudy (anonymous):

help please? @texaschic101

OpenStudy (anonymous):

someone please help me. im almost out of time.

OpenStudy (anonymous):

The answer is A: parallelogram. The points on the graph form a parallelogram.

OpenStudy (anonymous):

help please

OpenStudy (texaschic101):

you got more ?

OpenStudy (anonymous):

1 more.

OpenStudy (anonymous):

@texaschic101? please just 1 more.

OpenStudy (texaschic101):

what ya got

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

A triangle has a base of 11 inches and a height of 8 inches. What is the area of the triangle? A.121 square inches B.88 square inches C.64 square inches D.44 square inches im terrible with area

OpenStudy (anonymous):

Multiply the base by the height as if you were finding the area of a square. Then, divide that by 2 to find the area of a triangle, as a triangle is 1/2 of a square.

OpenStudy (texaschic101):

A = b * 1/2 * h A = 11 * 1/2(8) A = 11 * 4 A = 44